# Topology-Informed Graph Transformer

### v1

@@ 4 EXPERIMENTS
Dataset     To analyze the effectiveness of TIGT compared to other models in terms of expressive powers, we experiment on the Circular Skip Link(CSL) dataset (Murphy et al., 2019). CSL dataset is comprised of graphs that have different skip lengths $R\in\{2,3,4,5,6,9,11,12,13,16\}$ with 41 nodes that have the same features. Further, we utilize well-known graph-level benchmark datasets to evaluate proposed models compared to other models. We leverage five datasets from the ”Benchmarking GNN” studies: MNIST, CIFAR10, PATTERN, and CLUSTER, adopting the same experimental settings as prior research (Dwivedi et al., 2022). Additionally, we use two datasets from the ”Long-Range Graph Benchmark” (Dwivedi et al., 2023): Peptides-func and Peptides-struct. Lastly, to further verify the effectiveness of the proposed model on large datasets, we perform experiments on ZINC full dataset (Irwin et al., 2012), which is the full version of the ZINC dataset with 250K graphs and PCQM4Mv2 dataset (Hu et al., 2020) which is large-scale graph regression benchmark with 3.7M graphs. These benchmark encompass binary classification, multi-label classification, and regression tasks across a diverse range of domain characteristics. The detail of the aforementioned datasets are summarized in Appendix C.1.
Models     To evaluate the discriminative power of TIGT, we compare a set of previous researches related to expressive power of GNNs on CSL dataset such as Graph Transformers (GraphGPS (Rampášek et al., 2023), GRIT (Ma et al., 2023)) and other message-passing neural networks (GCN (Kipf & Welling, 2017), GIN, Relational Pooling GIN(RP-GIN) (Murphy et al., 2019), Cy2C-GNNs (Choi et al., 2023)). We compare our approach on well-known benchmark datasets to test graph-level test with the latest SOTA techniques, widely adopted MPNNs models, and various Graph Transformer-based studies: GRIT (Ma et al., 2023), GraphGPS (Rampášek et al., 2023)), GCN (Kipf & Welling, 2017), GIN (Xu et al., 2019), its variant with edge-features (Hu et al., 2020), GAT (Veličković et al., 2018), GatedGCN (Bresson & Laurent, 2018), GatedGCN-LSPE (Dwivedi et al., 2022), PNA (Corso et al., 2020), Graphormer (Ying et al., 2021), K-Subgraph SAT (Chen et al., 2022), EGT (Hussain et al., 2022), SAN (Kreuzer et al., 2021), Graphormer-URPE (Luo et al., 2022), Graphormer-GD (Zhang et al., 2023), DGN (Beaini et al., 2021), GSN  (Bouritsas et al., 2021), CIN (Bodnar et al., 2021b), CRaW1  (Tönshoff et al., 2023), and GIN-AK+ (Zhao et al., 2022).
@@
Performance on the CSL dataset     In order to test the expressive power of the proposed model and state-of-the-art Graph Transformers, we evaluated their performance on the synthetic dataset CSL. The test performance metrics are presented in Table 1. Our analysis found that TIGT, GPS with random-walk structural encoding (RWSE), and GPS with RWSE and Laplacian eigenvectors encodings (LapPE) outperformed other models. However, the recent state-of-the-art model, GRIT with Relative Random Walk Probabilities (RRWP), could not discriminate CSL class. Interestingly, TIGT demonstrated resilience in maintaining a near 100% performance rate, irrespective of the number of added Graph Transformer layers. This consistent performance can be attributed to TIGT’s unique Dual-path message-passing layer, which ceaselessly infuses topological information across various layers. Conversely, other models, which initially derive benefits from unique node attribution facilitated by positional encoding, showed signs of diminishing influence from this attribution as the number of layers grew. Additionally, we compared our findings with those of GAT and Cy2C-GNNs models. Consistent with previous studies (Choi et al., 2023), GAT was unable to perform the classification task on the CSL dataset effectively. In the case of the Cy2C-GNN model, while it demonstrated high accuracy in a single-layer configuration, similar to GPS, we observed a decline in classification performance as the number of layers increased.
@@
Results from benchmark datasets     First, we present the test performance on five datasets from Benchmarking GNNs (Dwivedi et al., 2022) in Table 2. The mean and standard deviation are reported over four runs using different random seeds. It is evident from the results that our model ranks either first or second in performance on three benchmark datasets: ZINC, MNIST, and CIFAR10. However, for the synthetic datasets, PATTERN, and CLUSTER, our performance is found to be inferior compared to recent state-of-the-art models but is on par with the GraphGPS model. Next, we further assess the effectiveness of our current model by evaluating its test performance on four datasets from the ”Long-Range Graph Benchmark” (Dwivedi et al., 2023), full ZINC dataset (Irwin et al., 2012), and the PCQM4Mv2 dataset (Hu et al., 2020). In the large datasets, the full version of the ZINC dataset and the PCQM4Mv2 dataset, TIGT consistently outperforms other models. In particular, on the PCQM4Mv2 dataset, our model demonstrated superior performance with fewer parameters compared to state-of-the-art models. In the ”Long-Range Graph Benchmark,” our model also present the second-highest performance compared to other models. Through all these experimental results, it is evident that by enhancing the discriminative power to differentiate isomorphisms of graphs, we can boost the predictive performances of Graph Transformers. This has enabled us to achieve competitive results in GNN research, surpassing even recent state-of-the-art model on several datasets. In a comparative analysis between Cy2C-GNN and TIGT, we observed a significant increase in performance across all datasets with TIGT. This indicates that the topological non-trivial features of graphs are well-reflected in TIGT, allowing for both a theoretical increase in expressive power and improved performance on benchmark datasets.
@@
Table 1: Results of graph classification obtained from CSL dataset (Murphy et al., 2019). Note that a bold method indicate the results obtained by ours. All results other than the bold method are cited from available results obtained from pre-existing publications. Note that mean $\pm$ standard deviation of 4 runs with different random seeds in our results.Highlighted are the top first, second, and third results.
@@
Table 2: Graph classification and regression results obtained from five benchmarks from (Dwivedi et al., 2022). Note that N/A indicate the methods which do not report test results on the given graph data set and a bold method indicate the results obtained by ours. All results other than the bold method are cited from available results obtained from pre-existing publications. Note that mean $\pm$ standard deviation of 4 runs with different random seeds in our results.Highlighted are the top first, second, and third results.
@@
Table 3: Graph-level task results obtained from two long-range graph benchmarks (Dwivedi et al., 2023) , ZINC-full dataset (Irwin et al., 2012) and PCQM4Mv2 (Hu et al., 2020). Note that N/A indicate the methods which do not report test results on the given graph data set and a bold method indicate the results obtained by ours. All results other than the bold method are cited from available results obtained from pre-existing publications. Note that mean $\pm$ standard deviation of 4 runs with different random seeds in our results. Highlighted are the top first, second, and third results.
@@ 5 CONCLUSION
In this paper, we introduced TIGT, a novel Graph Transformer designed to enhance the predictive performance and expressive power of Graph Transformers. This enhancement is achieved by incorporating a topological positional embedding layer, a dual-path message passing layer, a global attention layer, and a graph information layer. Notably, our topological positional embedding layer is learnable and leverages MPNNs. It integrates universal covers drawn from the original graph structure and a modified structure enriched with cyclic subgraphs. This integration aids in detecting isomorphism classes. Throughout its architecture, TIGT encodes cyclic subgraphs at each layer using the dual-path message passing mechanism, ensuring that expressive power is maintained as layer depth increases. Despite a modest rise in complexity, TIGT showcases superior performance in experiments on the CSL dataset, surpassing the expressive capabilities of previous GNNs and Graph Transformers. Additionally, both mathematical justifications and empirical evaluations underscore our model’s competitive advantage over contemporary Graph Transformers across diverse benchmark datasets.
While TIGT can be successfully applied to graph-level tasks, there remain avenues for future exploration. Firstly, the computational complexity is limited to $O(N^{2}+N_{E}+N_{C})$ with the number of node $N$, the number of edge $N_{E}$ and the number of edge in cyclic subgraphs $N_{C}$. Especially, due to the implementation of global attention in the Transformer, computational complexity poses challenges that we are keen to address in subsequent research. Moreover, beyond the realm of graph-level tasks, there is potential to broaden the application of TIGT into areas like node classification and link prediction. Integrating the topological characteristics inherent in TIGT with these domains might uncover more profound insights and elevate predictive accuracy.
@@ Appendix B Abalation study
To understand the significance of each component in our deep learning model, we performed multiple ablation studies using the ZINC dataset (Dwivedi et al., 2022). The results are presented in Table 4. The influence of the graph information and the topological positional embedding layer is relatively marginal compared to other layers. The choice of weight-sharing within the topological positional embedding layer, as well as the selection between the hyperbolic tangent and ReLU activation functions, play a significant role in the model’s performance. Likewise, opting for Single-path MPNNs, excluding the adjacency matrix instead of the proposed Dual-path in each TIGT layer, results in a considerable performance drop. Within the graph information layer, it’s evident that employing a sum-based readout function, akin to graph pooling, is crucial for extracting comprehensive graph information and ensuring optimal results. Additionally, we experimented with applying the Performer, which utilizes a kernel trick to replace the quadratic complexity of the transformer’s global attention with linear complexity, in our TIGT model. However, we found that this resulted in performance similar to models that did not use global attention. This suggests that TIGT may require further research to address the issue of quadratic complexity effectively. In a similar setting, we conducted experiments with Cy2C-GNN, which has fewer parameters (114,433) compared to TIGT, and observed poorer performance. We also tested a larger version of Cy2C-GNN, named Cy2C-GNN(Large), with 1,766,401 parameters—approximately three times more than TIGT’s 539,873—only to find that this resulted in a worse mean absolute error (MAE).
@@ C.1 Datasets
C.2 Hyperparameters
@@
C.3 Implementation detail of Experiment on CSL dataset

### v2

@@ 4 EXPERIMENTS
With the theoretical analysis of TIGT in place, we present the empirical effectiveness of TIGT in discerning various properties of graph datasets.
Dataset     To analyze the effectiveness of TIGT compared to other models in terms of expressive powers, we experiment on the Circular Skip Link(CSL) dataset (Murphy et al., 2019). CSL dataset comprises of graphs that have different skip lengths $R\in\{2,3,4,5,6,9,11,12,13,16\}$ with 41 nodes that have the same features. Further, we utilize well-known graph-level benchmark datasets to evaluate proposed models compared to other models. We leverage five datasets from the ”Benchmarking GNN” studies: MNIST, CIFAR10, PATTERN, and CLUSTER, adopting the same experimental settings as prior research (Dwivedi et al., 2022). Additionally, we use two datasets from the ”Long-Range Graph Benchmark” (Dwivedi et al., 2023): Peptides-func and Peptides-struct. Lastly, to further verify the effectiveness of the proposed model on large datasets, we perform experiments on ZINC full dataset (Irwin et al., 2012), which is the full version of the ZINC dataset with 250K graphs and PCQM4Mv2 dataset (Hu et al., 2020) which is large-scale graph regression benchmark with 3.7M graphs. These benchmark encompass binary classification, multi-label classification, and regression tasks across a diverse range of domain characteristics. The details of the aforementioned datasets are summarized in Appendix C.1.
Models     To evaluate the discriminative power of TIGT, we compare a set of previous researches related to expressive power of GNNs on CSL dataset such as Graph Transformers (GraphGPS (Rampášek et al., 2023), GRIT (Ma et al., 2023)) and other message-passing neural networks (GCN (Kipf & Welling, 2017), GIN, Relational Pooling GIN(RP-GIN) (Murphy et al., 2019), Cy2C-GNNs (Choi et al., 2023)). We compare our approach on well-known benchmark datasets to test graph-level tasks with the latest SOTA techniques, widely adopted MPNNs models, and various Graph Transformer-based studies: GRIT (Ma et al., 2023), GraphGPS (Rampášek et al., 2023)), GCN (Kipf & Welling, 2017), GIN (Xu et al., 2019), its variant with edge-features (Hu et al., 2020), GAT (Veličković et al., 2018), GatedGCN (Bresson & Laurent, 2018), GatedGCN-LSPE (Dwivedi et al., 2022), PNA (Corso et al., 2020), Graphormer (Ying et al., 2021), K-Subgraph SAT (Chen et al., 2022), EGT (Hussain et al., 2022), SAN (Kreuzer et al., 2021), Graphormer-URPE (Luo et al., 2022), Graphormer-GD (Zhang et al., 2023), DGN (Beaini et al., 2021), GSN  (Bouritsas et al., 2021), CIN (Bodnar et al., 2021b), CRaW1  (Tönshoff et al., 2023), and GIN-AK+ (Zhao et al., 2022).
@@
Performance on the CSL dataset     In order to test the expressive power of the proposed model and state-of-the-art Graph Transformers, we evaluated their performance on the synthetic dataset CSL. The test performance metrics are presented in Table 1. Our analysis found that TIGT, GPS with random-walk structural encoding (RWSE), and GPS with RWSE and Laplacian eigenvectors encodings (LapPE) outperformed other models. However, the recent state-of-the-art model, GRIT with Relative Random Walk Probabilities (RRWP), could not distinguish the CSL classes. Interestingly, TIGT demonstrated resilience in maintaining a near 100% performance rate, irrespective of the number of added Graph Transformer layers. This consistent performance can be attributed to TIGT’s unique Dual-path message-passing layer, which ceaselessly infuses topological information across various layers. Conversely, other models, which initially derive benefits from unique node attributes facilitated by positional encoding, showed signs of diminishing influence from this attribution as the number of layers grew. Additionally, we compared our findings with those of GAT and Cy2C-GNN models. Consistent with previous studies (Choi et al., 2023), GAT was unable to perform the classification task on the CSL dataset effectively. In the case of the Cy2C-GNN model, while it demonstrated high accuracy in a single-layer configuration, similar to GPS, we observed a decline in classification performance as the number of layers increased.
@@
Results from benchmark datasets     First, we present the test performance on five datasets from Benchmarking GNNs (Dwivedi et al., 2022) in Table 2. The mean and standard deviations are reported over four runs using different random seeds. It is evident from the results that our model ranks either first or second in performance on three benchmark datasets: ZINC, MNIST, and CIFAR10. However, for the synthetic datasets, PATTERN, and CLUSTER, our performance is found to be inferior compared to recent state-of-the-art models but is on par with the GraphGPS model. Next, we further assess the effectiveness of our current model by evaluating its test performance on four datasets from the ”Long-Range Graph Benchmark” (Dwivedi et al., 2023), full ZINC dataset (Irwin et al., 2012), and the PCQM4Mv2 dataset (Hu et al., 2020). In large datasets, such as the full version of the ZINC dataset and the PCQM4Mv2 dataset, TIGT consistently outperforms other models. In particular, on the PCQM4Mv2 dataset, our model demonstrated superior performance with fewer parameters compared to state-of-the-art models. In the ”Long-Range Graph Benchmark,” our model also presents the second-highest performance compared to other models. Through all these experimental results, it is evident that by enhancing the discriminative power to differentiate isomorphisms of graphs, we can boost the predictive performance of Graph Transformers. This has enabled us to achieve competitive results in GNN research, surpassing even recent state-of-the-art model on several datasets. In a comparative analysis between Cy2C-GNN and TIGT, we observed a significant increase in performance across all datasets with TIGT. This indicates that the topological non-trivial features of graphs are well-reflected in TIGT, allowing for both a theoretical increase in expressive power and improved performance on benchmark datasets.
@@
Table 1: Results of graph classification obtained from CSL dataset (Murphy et al., 2019). Note that the methods written in bolded texts indicate the results obtained from our implementations. All results other than the bolded methods are cited from available results obtained from pre-existing publications. Our results are written as “mean $\pm$ standard deviation”, which are obtained from 4 runs with different random seeds. Highlighted are the top first, second, and third results.
@@
Table 2: Graph classification and regression results obtained from five benchmarks from (Dwivedi et al., 2022). Note that N/A indicate the methods which do not report test results on the given graph data set. The methods written in bolded texts indicate the results obtained from our implementations. All results other than the bolded methods are cited from available results obtained from pre-existing publications. Our results are written as “mean $\pm$ standard deviation”, which are obtained from 4 runs with different random seeds. Highlighted are the top first, second, and third results.
@@
Table 3: Graph-level task results obtained from two long-range graph benchmarks (Dwivedi et al., 2023) , ZINC-full dataset (Irwin et al., 2012) and PCQM4Mv2 (Hu et al., 2020). Note that N/A indicate the methods which do not report test results on the given graph data set. The methods written in bolded texts indicate the results obtained from our implementations. All results other than the bolded methods are cited from available results obtained from pre-existing publications. Our results are written as “mean $\pm$ standard deviation”, which are obtained from 4 runs with different random seeds. Highlighted are the top first, second, and third results.
@@ 5 CONCLUSION
In this paper, we introduce TIGT, a novel Graph Transformer designed to enhance the predictive performance and expressive power of Graph Transformers. This enhancement is achieved by incorporating a topological positional embedding layer, a dual-path message passing layer, a global attention layer, and a graph information layer. Notably, our topological positional embedding layer is learnable and leverages MPNNs. It integrates universal covers drawn from the original graph structure and a modified structure enriched with cyclic subgraphs. This integration aids in detecting isomorphism classes of graphs. Throughout its architecture, TIGT encodes cyclic subgraphs at each layer using the dual-path message passing mechanism, ensuring its expressive power to be maintained as layer depth increases. Despite a modest rise in complexity, TIGT showcases superior performance in experiments on the CSL dataset, surpassing the expressive capabilities of previous GNNs and Graph Transformers. Additionally, both mathematical justifications and empirical evaluations underscore our model’s competitive advantage over contemporary Graph Transformers across diverse benchmark datasets.
While TIGT can be successfully applied to graph-level tasks, there remain avenues for future exploration. Firstly, the computational complexity is limited to $O(N^{2}+N_{E}+N_{C})$ with the number of nodes $N$, the number of edges $N_{E}$ and the number of edge in cyclic subgraphs $N_{C}$. Especially, due to the implementation of global attention in the Transformer, computational complexity poses challenges that we are keen to address in subsequent research. Moreover, beyond the realm of graph-level tasks, there is potential to broaden the application of TIGT into areas like node classification and link prediction. Integrating the topological characteristics inherent in TIGT with these domains might uncover more profound insights and enhance predictive accuracy.
@@ Appendix B Abalation study
To understand the significance of each component in our deep learning model, we performed multiple ablation studies using the ZINC dataset (Dwivedi et al., 2022). The results are presented in Table 4. The influence of the graph information and the topological positional embedding layer is relatively marginal compared to other layers. The choice of weight-sharing within the topological positional embedding layer, as well as the selection between the hyperbolic tangent and ReLU activation functions, play a significant role in the model’s performance. Likewise, opting for Single-path MPNNs, excluding the adjacency matrix instead of the proposed Dual-path in each TIGT layer, results in a considerable performance drop. Within the graph information layer, it’s evident that employing a sum-based readout function, akin to graph pooling, is crucial for extracting comprehensive graph information and ensuring optimal results. Additionally, we experimented with applying the Performer, which utilizes a kernel trick to replace the quadratic complexity of the transformer’s global attention with linear complexity, in our TIGT model. However, we found that this resulted in performance similar to models that did not use global attention. This suggests further research on effectively addressing the issue of quadratic complexity inherent in TIGT. In a similar setting, we conducted experiments with Cy2C-GNN, which has fewer parameters (114,433) compared to TIGT, and observed poorer performance. We also tested a larger version of Cy2C-GNN, named Cy2C-GNN(Large), with 1,766,401 parameters—approximately three times more than TIGT’s 539,873—only to find that this resulted in a worse mean absolute error (MAE).
@@ C.1 Datasets
C.2 Additional notes on model design
TIGT includes batch normalization layer, proposed from  (Ioffe & Szegedy, 2015), and residual connection, as outlined in  (He et al., 2016).
C.3 Hyperparameters
@@
C.4 Implementation detail of Experiment on CSL dataset